INFRARED SPECTROSCOPY OF METAL LIGAND AND METAL OXIDE LIGAND COMPLEXES

\underline{ALLEN M. RICKS} and MICHAEL A. DUNCAN, Department of Chemistry, University of Georgia, Athens, Ga.

Metal ligand and metal oxide ligand complexes are generated in a laser vaporization/supersonic expansion cluster source and are mass selected and studied using infrared spectroscopy in the 600-4000 cm^{−1}region using a tunable infrared OPO. Several different systems are investigated and some show evidence for intracluster reactions to form metal oxide species or coupling of coordinating ligands. In particular the V(CO_{2})_{n}^{+} system shows evidence for the onset of in intracluster reaction at n=7 as evidinced by the sudden appearance of a band at 1800 cm^{−1}. DFT calculations are performed in support of this work and several possible interpretations of the experimental data will be presented.